# Comparison of Nested Geometry Treatments within GPU-Based Monte Carlo Neutron Transport Simulations of Fission Reactors

### v1

@@ 6 Performance testing
Single node performance testing trials, which were run with only 10 inactive and 10 active cycles, were not sufficient to produce converged $k_{\mathrm{eff}}$ results. To confirm that all four tracking methods produce the same solution, a final trial was performed for each method on 50 nodes of Frontier with a workload of $2.5\times 10^{5}$ histories/cycle/GCD, for a total of $10^{8}$ histories per cycle. For each trial, 200 inactive cycles and 400 active cycles were run. A lower workload was necessary to complete all 600 cycles within Frontier’s walltime limit. Converged $k_{\mathrm{eff}}$ values agreed closely, as shown in Table 5. Neutron flux results for all 4 methods also matched expectations. Figure 9 shows the converged flux for the SP method trial. This figure clearly shows the depression on the flux within the borosilicate glass rods, which act as neutron absorbers.
@@ 7 Demonstration problem
In Section 6, the SP method consistently provided the best performance, but all three methods achieved tracking rates reasonably close to that of RTK. From a software engineering perspective, when considering multiple approaches, the trade-off between performance and other factors such as code simplicity and maintainability must be assessed. The ST method achieves over 90% of the SP method tracking rate in all cases shown in Table 2 and has the added benefit of not requiring an additional hexagonal array tracker. This is advantageous for hexagonal-array-based reactors, which cannot be represented using RTK, and would otherwise require the development of a GPU-based hexagonal array tracker for use with the SP method. To demonstrate this benefit, the ST method was used to perform a $k_{\mathrm{eff}}$ calculation on the Empire microreactor benchmark problem, which consists of nested hexagonal arrays as shown in Figure 10. Microreactors, even smaller than SMRs, are designed to be factory assembled and easily transported in order to supply power for remote areas, disaster relief, and space applications.
@@
The problem was run on 100 nodes on Frontier using all eight GCD s on each node, with a workload of $10^{6}$ histories/cycle/GCD, for a total of $8\times 10^{8}$ histories/cycle. To achieve converged results, 120 inactive cycles and 120 active cycles where performed. The flux and fission source were tallied on an extremely fine rectilinear mesh with 3350 mesh elements in $x$ and $y$ (0.05 cm resolution) and a single mesh element in $z$ spanning $\pm$0.25 cm on the midplane (11,222,500 mesh volume elements total). A converged $k_{\mathrm{eff}}$ of 1.026777 $\pm$ 0.000004 was obtained. Converged flux results are shown in Figure 11. This figure shows the depression in the flux around the neutron absorbers and interstitial spaces between assemblies where no fuel is present. Figure 12 shows the converged fission source, which is highly resolved within each fuel rod. The highest fission source intensity occurs around the edges of fuel rods. For rods in the outer parts of the assembly, the fission source is observed to be highest on the inward-facing edge.
Figure 11: Converged neutron flux distribution on the midplane of the Empire problem using the ST method, obtained on 100 nodes of Frontier with $10^{6}$ histories/cycle/GCD, with 120 inactive and 240 active cycles. Statistical errors in the flux on this slice are all less than 0.7% within the core region.
@@
Table 6 shows tracking rates, as well as GPU vs. CPU comparison results, in the same format as that used in Table 4. For inactive cycles, the tracking rate of $5.29\times 10^{4}$ histories/s/GCD is 40.4% of the ST method tracking rate for the NuScale problem on Frontier. For active cycles, the tracking rate of $1.02\times 10^{4}$ histories/s/GCD is only 14.7% of the ST method tracking rate for the NuScale problem on Frontier. This is reasonable because the mesh used for the Empire problem is 26.4$\times$ larger than the mesh used for the NuScale problem and is much larger than meshes used for most practical applications. With respect to CPU execution, a modest 1.72$\times$ speedup was achieved for inactive cycles; however, for active cycles, a 2.94$\times$ slowdown was observed. The stark difference between inactive and active cycle performance can be attributed to tallying operations, which only occur during active cycles. Tally operations accounted for 61% of the total GPU runtime (i.e., runtime spend in both inactive and active cycles) as shown in Figure 13. These results indicate that more work is necessary to improve the performance of tally operations on the GPU.
Table 6: GPU to CPU performance comparison for the Empire problem on Frontier, with $10^{6}$ histories/cycle/GCD. Speedup refers to the reduction in time-to-solution from using all GPUs on a node compared to all CPUs.
@@
(neutron/s/GCD) per GPU
Inactive $5.29\times 10^{4}$ 12.0 1.72
Active $1.02\times 10^{4}$ 2.37 0.34
Figure 13: Percentage of total GPU runtime (both inactive and active cycles) spent performing geometry, physics, tally, and other operations for the Empire problem on Frontier. Cross section calculations (which typically dominate runtime) fall within the “physics” category.
@@ 8 Conclusion
In this work, three methods for GPU-based neutron tracking within multi-universe geometries where tested and compared to RTK. For the NuScale problem, the DP, SP, and ST methods spent 2.45$\times$, 1.50$\times$, and 2.02$\times$ more time in tracking kernels in inactive cycles on Frontier compared to RTK. However, although inactive cycles on this fresh fuel problem represent the maximum time spent performing tracking operations, these methods still achieved 77.8%, 91.2%, and 83.4% of the RTK tracking rate. It is concluded that any of these three methods—which unlike RTK can handle arbitrarily nested configurations—are suitable for production-level use, incurring only a modest performance penalty.
The SP method outperformed the DP and ST methods in nearly all cases. However, the ST method achieved over 90% of the SP method’s tracking rate in all cases, and unlike the SP method, it requires only a single CSG tracker. The flexibility of this approach was demonstrated with the Empire microreactor, a hexagonal array-based reactor which cannot be represented using RTK, and would require an additional GPU-based hexagonal array tracker to be written in order to use the SP method. The ST method was used to produce converged $k_{\mathrm{eff}}$, neutron flux, and fission source results, which matched expectations. A modest speedup of 1.72$\times$ over CPU execution on Frontier was observed during inactive cycles. Additional work is necessary outside the neutron tracking portions of Shift to remediate the a 2.94$\times$ slowdown observed for active cycles, understanding that the mesh tally used for this problem is far larger than mesh tallies used for typical applications. This work will facilitate GPU-based MC transport for reactor problems with geometry complexity beyond rectilinear arrays, as well as other non-reactor radiation transport problems with complex nested geometries.

### v2

@@ 6.5 Verification
Single node performance testing trials, which were run with only 10 inactive and 10 active cycles, were not sufficient to produce converged $k_{\mathrm{eff}}$ results. To verify that all four tracking methods produce the same solution, a final trial was performed for each method on 50 nodes of Frontier with a workload of $2.5\times 10^{5}$ histories/cycle/GCD, for a total of $10^{8}$ histories per cycle. For each trial, 200 inactive cycles and 400 active cycles were run. A lower workload was necessary to complete all 600 cycles within Frontier’s walltime limit. Converged $k_{\mathrm{eff}}$ values agreed closely, as shown in Table 6. Neutron flux results for all 4 methods also matched expectations. Figure 9 shows the converged flux for the SP method trial. This figure clearly shows the depression on the flux within the borosilicate glass rods, which act as neutron absorbers.
@@ 7 Demonstration problem
In Section 6.1, the SP method consistently provided the best performance, but all three methods achieved tracking rates reasonably close to that of RTK. From a software engineering perspective, when considering multiple approaches, the trade-off between performance and other factors such as code simplicity and maintainability must be assessed. The ST method achieves over 90% of the SP method tracking rate in all cases shown in Table 2 and has the added benefit of not requiring an additional hexagonal array tracker. This is advantageous for hexagonal-array-based reactors, which cannot be represented using RTK, and would otherwise require the development of a GPU-based hexagonal array tracker for use with the SP method. To demonstrate this benefit, the ST method was used to perform a $k_{\mathrm{eff}}$ calculation on the Empire microreactor benchmark problem that consists of nested hexagonal arrays, as shown in Figure 10. Microreactors, even smaller than SMRs, are designed to be factory assembled and easily transported in order to supply power for remote areas, disaster relief, and space applications.
@@
A Cartesian mesh tally would require a prohibitively large number of elements in order to conform to the hexagonal pins and assemblies found in the Empire geometry, and Shift does not yet support hexagonal mesh tallies on the GPU. Thus, cell-based tallies were used to tally the flux and fission source density within each fuel pin. For high-resolution results, each fuel pin was subdivided into 360 cells, with 3 radial divisions, 4 circumferential divisions, and 30 axial divisions. This resulted in a total of 388,800 cell tallies, similar to the number of elements in the Cartesian mesh tally used for the NuScale SMR problem.
The problem was run on 100 Frontier nodes using all eight GCDs on each node, with a workload of $10^{6}$ histories/cycle/GCD, for a total of $8\times 10^{8}$ histories/cycle. To achieve converged results, 120 inactive cycles and 120 active cycles were performed. For comparison, CPU results were obtained on a single node of Frontier with $2\times 10^{6}$ histories/cycle for 120 inactive cycles and 120 active cycles.
Converged flux results are shown in Figure 11. As expected, the highest flux occurs in an annular region around the center of the reactor, i.e., the region where leakage to the outside of the reactor and inner void region are minimized. Figure 12 shows the converged fission source. As expected, the highest fission source density occurs around the edges of fuel rods. For rods in the outer parts of assemblies, the fission source density is observed to be highest on the inward-facing edge. A converged $k_{\mathrm{eff}}$ of 1.026772 $\pm$ 0.000003 was obtained.
Figure 11: Converged neutron flux distribution on the midplane of the Empire problem using the ST method, obtained on 100 nodes of Frontier with $10^{6}$ histories/cycle/GCD, with 120 inactive and 120 active cycles. Statistical uncertainties on this slice are all less than 0.10%.
Figure 12: Converged fission source distribution on the midplane of the Empire problem using the ST method, obtained on 100 nodes of Frontier with $10^{6}$ histories/cycle/GCD, with 120 inactive and 120 active cycles. Statistical uncertainties on this slice are all less than 0.25%.
@@
Table 7: GPU to CPU performance comparison for the Empire problem on Frontier, with $10^{6}$ histories/cycle/GCD. Speedup is defined as the ratio of the GPU tracking rate using all GPUs on a single node to the CPU tracking rate using all CPU cores on a single node.
@@
(histories/s/GCD) per GCD
Inactive $4.74\times 10^{4}$ 15.9 2.27
Active $4.33\times 10^{4}$ 14.8 2.12
@@ 8 Conclusion
[…]